GIven that log36=m, and log65=n express log310 in terms of m and n



1
Expert's answer
2021-10-25T03:33:36-0400

GIven, "log_36"=m, and "log_65" =n.

"log_36=m\\
\frac{log6}{log3}=m\space--------(1)\\
\frac{log(2×6)}{log3}=m\\
\frac{log2+log3}{log3}=m\\
\frac{log2}{log3}+\frac{log3}{log3}=m\\
\frac{log2}{log3}=m-1\\
And,\\
log_65=n\\
\frac{log5}{log6}=n\space---------(2)\\
\text{By multiple (1) and (2), we get}\\
\frac{log6}{log3}\frac{log5}{log6}=\frac{log5}{log3}=mn\space ----(3)\\
Then,\\
log_310=\frac{log10}{log3}\\
=\frac{log(2×5)}{log3}\\
=\frac{log2+log5}{log3}\\
=\frac{log2}{log3}+\frac{log5}{log3}\\
=m-1+mn\space \space[from \space(3)]\\
=m(1+n)-n"